OpenStudy (ilovehorses14):

At a video arcade, the action video games cost $22, which is 110% of last week’s price. What was the price last week? A. $15 B. $18 C. $20 D. $23

OpenStudy (anonymous):

The easiest way to do this is to take each given answer price and add 10% to it. So: A. 15+1.5=16.5 B. 18+1.8=19.8 C. 20+2=22 D. 23+2.3=25.3 Now it's easy to see that the cost the previous week was $20.

OpenStudy (ilovehorses14):

Mrs. Manning’s current savings is 142% of what she had saved last month. If she has $923 in her account this month. What was her last month’s savings? A. $600 B. $623 C. $650 D. $745